OpenStudy (keyonsay):

PLEASE HELP!!!!!! Simplify the following complex fraction. 0.15/ 1.72- 1/2

OpenStudy (anonymous):

without calculator..(.15*100)/(1.72*100)-1/2 =15/172-1/2 =(15*2-172)/(172*2) i think u can simplify after this..the ans will be negative.

OpenStudy (keyonsay):

thank you :D

OpenStudy (anonymous):

you're welcome:)